OpenStudy (anonymous):

Can someone explain how to solve this step by step? -12=p-5p I think if I added terms it would be -12=-4p?

OpenStudy (anonymous):

-12=-4p 12=4p p=3

OpenStudy (dumbcow):

good so far :)

OpenStudy (anonymous):

So, you divide a negative # by a negative # and it is...positive?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

yes!!!!

OpenStudy (dumbcow):

\[\frac{-4}{-4}p = \frac{-12}{-4}\] \[p=3\]

OpenStudy (anonymous):

Ok. So if it was: -12=p-5p p=? That's what I'm confused about.

OpenStudy (dumbcow):

thats the exact same question you just solved?? p=3

OpenStudy (anonymous):

No, theres a 5.

OpenStudy (dumbcow):

ahh yes but you did the 1st step already p-5p = -4p

OpenStudy (anonymous):

Ok. So... -12+p-5p P is positive 3.

OpenStudy (dumbcow):

yeah, check it yourself plug in p=3 and see what you get

OpenStudy (anonymous):

Okay...... It is positive 3 because of the negative 4 that is divided into -12?

OpenStudy (dumbcow):

correct

OpenStudy (dumbcow):

and the reason we divided -4 was because it was the coefficient of p

OpenStudy (anonymous):

Yes, thank you! xD

OpenStudy (dumbcow):

no problem